OpenStudy (anonymous):

integral of (sinx)/(1-sin^2x) dx

OpenStudy (campbell_st):

use the trig identity sin^2 + cos^2 = 1 then you are integrating sin(x)/cos^2(x) = tan(x)sec(x)

sam (.sam.):

yes then \[\text{substitute }u=\sec (x)\text{ and }d u=\tan (x) \sec (x)d x\] \[\text{}=\int\limits 1 \, du\] \[=u+c\] \[\begin{array}{l} \text{Substitute back for }u=\sec (x): \\ \text{}=\sec (x)+\text{c} \\\end{array}\]

OpenStudy (anonymous):

thamks